Residual Income Model and Nonconstant Growth When a stock is going through a period of nonconstant growth for T periods, followed by constant growth forever, the residual income model can be modified as follows:

Where

Al’s Infrared Sandwich Company had a book value of $12.95 at the beginning of the year, and the earnings per share for the past year were $3.41. Molly Miller, a research analyst at Miller, Moore & Associates, estimates that the book value and earnings per share will grow at 12.5 and 11 percent per year for the next four years, respectively. After four years, the growth rate is expected to be 6 percent. Molly believes the required return for the company is 8.2 percent. What is the value per share for Al’s Infrared Sandwich Company?
Answer #1

Book Value

Book value refers to the value of an asset or investment that is written in the books of accounts of an organization. It serves as a very important role in the computation of profit or loss caused due to sale of the asset.

The values for year-end can be estimated as:

Book value = $12.95 * (1.1250) = $14.57

EPS = $3.41 * (1.11) = $3.79

 

Where,

The book value for the first year can be calculated by using the following approach:

B2 – B1 = B1 (1 + g) – B1 = B1 + B1g – B1 = B1g

In order to compute the book value of the other years use the following approach:

Therefore, the book value of the years computed using the above approach is